# Extracting transport coefficients from local ground-state currents

### v1

@@ IV Conclusion and Outlook
Owing to its broad applicability, our method can be utilized for a wide range of gapped many-body systems, such as fractional Chern insulators that exhibit intrinsic topological order; see Appendix D. This approach can also be utilized for mixed states, allowing for the investigation of transport coefficients at finite temperatures. Furthermore, our protocol extends to other (potentially multi-point) correlators, involving (local) densities and currents, thereby giving access to response functions hardly accessible in traditional solid-state experiments.
Acknowledgements.
Appendix A Oscillation frequency of $\mathcal{C}(\vec{r},t)$
To justify our statement that the oscillations of the current-current correlator
@@ Appendix A Oscillation frequency of $\mathcal{C}(\vec{r},t)$
In principle, the expression in Eq. (27) encompasses multiple energy gaps, resulting in oscillations at various frequencies. Nevertheless, the dominant contribution is expected to arise from transitions between the ground band and the first excited band, thereby producing a single characteristic frequency, $\omega_{0}$. This behavior is exemplified in the Landau-level limit, where selection rules restrict transitions from the lowest Landau level (n=0n\!=\!0) to the next (n=1n\!=\!1). Additional selection rules may emerge from specific symmetries, as seen in the Hofstadter model, which exhibits chiral symmetry [5].
We now explicitly demonstrate this argument through a numerical study of the Hofstadter model. We start by comparing the characteristic oscillation frequency $\omega_{0}$, extracted from a fit of the correlator $\mathcal{C}(\vec{r},t)$ using Eq. (6), to the cyclotron gap $\Omega_{c}$. The latter quantity corresponds to the bandgap to the first excited band in the Hofstadter model, and it reaches the value $\Omega_{c}^{\rm cont}=4\pi\alpha$ in the low-flux (continuum) limit. The results shown in Fig. 4 show a good agreement $\omega_{0}\!\approx\!\Omega_{c}$ within a broad flux window, $\alpha\lesssim 0.2$, hence validating the single-frequency ansatz in Eq. (6) in this range. Furthermore, the identification of the cyclotron bandgap as the relevant frequency indicates that the parameter $\omega_{0}$ can be determined independently through spectroscopy.
We also show the parameter $\omega_{0}$ as extracted from the coefficients $c_{m}$; see Eq. (12). This approach builds on the short-time approximation inherent to the expansion in Eq. (11), and agrees with the single-frequency ansatz [Eq. (6)] throughout the entire flux range.
Appendix B Locality of the Chern marker
To confirm the local nature of the Chern marker, we consider a variant of the Hofstadter model on an elongated rectangle of size $27\times 13$ with flux $\alpha=\pm\nicefrac{{1}}{{5}}$ in the left (right) half, respectively, separated by a single column of local potential $J\sum_{y=1}^{13}\hat{n}_{(14,y)}$ in the middle. By numerically evaluating $\mathcal{C}(\vec{r},t)$ [Eq. (4)] and extracting the local Hall response $\sigma_{\rm H}(\vec{r})$ from a fit with Eq. (6), we find that the local Chern marker ${\sf Ch}(\vec{r})=2\pi\sigma_{\rm H}(\vec{r})$ takes values close to +1+1 (1-1) in the left (right) half of the system according to the sign of the magnetic flux, see Fig. 5, thus directly confirming its local nature.
Figure 5: Density and local Chern marker as obtained from the fit in Eq. (6) for a Hofstadter model with opposite flux $\pm\alpha=\nicefrac{{1}}{{5}}$ per plaquette in the left and right half separated by a single column of a local potential. As expected, we find that the local Chern marker ${\sf Ch}(\vec{r})$ changes sign as the reference site $\vec{r}$ is moved from one region to the other.
Appendix C Including higher coefficients $c_{m}$
In order to improve the accuracy of our protocol, it is natural to include higher coefficients in the expansion of the correlator $\mathcal{C}(\vec{r},t)$ in Eq. (9),
@@ Appendix C Including higher coefficients $c_{m}$
Appendix D Laughlin state of hard-core bosons
We validate the applicability of our scheme to strongly correlated matter by applying it to the bosonic Laughlin state [31]. Specifically, we consider the Hofstadter model discussed in the main text, and add repulsive Hubbard interactions of strength $U>0$,
For simplicity, we consider the hard-core bosonic limit ($U/J\to\infty$), where the exact expressions for the commutators in Eq. (16) of the main text remain unchanged. At magnetic filling factor $\nu=\nicefrac{{1}}{{2}}$, it is well-established that this model hosts a lattice analog of the corresponding Laughlin state [32].
We limit our simulations to a system of three particles on $5\times 5$ sites at flux $\alpha=\nicefrac{{1}}{{4}}$ per plaquette, and perform exact diagonalization of the Hamiltonian to find the ground state of the model. We perform exact time-evolution to obtain the current-current correlator $\mathcal{C}(\vec{r},t)$ for a reference site $\vec{r}$ located at the center of the system.
Figure 6 shows a fit to this data, using the ansatz in Eq. (6). From this fit, we extract the local Chern marker ${\sf Ch}(\vec{r})=2\pi\sigma_{\rm H}(\vec{r})=0.34$. We attribute the deviation from the expected many-body Chern number ${\sf Ch}=\nicefrac{{1}}{{2}}$ to significant band-dispersion and finite-size effects.
We observe that the characteristic frequency $\omega_{0}$ extracted from the fit approximately matches the cyclotron gap. We attribute this to residual Landau-level selection rules, which are expected to become exact in the low-flux limit of the Hofstadter model; see also Ref. [33].
Furthermore, we apply the protocol described in the main text using three parameters ($c_{0},c_{1},c_{2}$), which yields ${\sf Ch}(\vec{r})\!=\!0.47$. Considering five parameters ($c_{0},\ldots,c_{4}$), we find ${\sf Ch}(\vec{r})\!=\!0.40$; see Fig. 6.
Our findings indicate that despite the simplifying assumptions and strong correlations within the Laughlin state, our approach is capable of effectively extracting a local topological marker with a satisfactory level of accuracy. This is due to the finite correlation length present in the Laughlin state, originating from the many-body bulk gap. This key property enables the establishment of a local Chern marker in strongly-correlated topological states, as similarly demonstrated by the local Streda response [9].
Figure 6: Time-dependent current-current correlations $\mathcal{C}(\vec{r},t)$ for three hard-core bosons on $5\times 5$ sites with $\alpha=\nicefrac{{1}}{{4}}$. The fit with Eq. (6) as well as the experimental protocol give estimates for the local Chern marker consistent with the many-body Chern number ${\sf Ch}=\nicefrac{{1}}{{2}}$ of the Laughlin state at $\nu=\nicefrac{{1}}{{2}}$ up to finite-size effects.

### v2

@@ Discussion
Owing to its broad applicability, our approach can also be utilized for mixed states, allowing for the investigation of transport coefficients at finite temperatures. Furthermore, our protocol extends to other (potentially multi-point) correlators, involving (local) densities and currents, thereby giving access to response functions hardly accessible in traditional solid-state experiments.
Methods
Oscillation frequency of $\mathcal{C}(\vec{r},t)$. To justify our statement that the oscillations of the current-current correlator
@@ Methods
In principle, the expression in Eq. (28) encompasses multiple energy gaps, resulting in oscillations at various frequencies. Nevertheless, the dominant contribution is expected to arise from transitions between the ground band and the first excited band, thereby producing a single characteristic frequency, $\omega_{0}$. This behavior is exemplified in the Landau-level limit, where selection rules restrict transitions from the lowest Landau level (n=0n\!=\!0) to the next (n=1n\!=\!1). Additional selection rules may emerge from specific symmetries, as seen in the Harper-Hofstadter model, which exhibits chiral symmetry [24].
The validity of this argument can be confirmed by comparing the characteristic oscillation frequency $\omega_{0}$, extracted from a fit of the correlator $\mathcal{C}(\vec{r},t)$ using Eq. (6), to the cyclotron gap $\Omega_{c}$, i.e. the bandgap to the first excited band in the Harper-Hofstadter model, and it reaches the value $\Omega_{c}^{\rm cont}=4\pi\alpha$ in the low-flux (continuum) limit; see Fig. 4 in the Results section.
Locality of the Chern marker. To confirm the local nature of the Chern marker, we consider a variant of the Harper-Hofstadter model on an elongated rectangle of size $27\times 13$ with flux $\alpha=\pm\nicefrac{{1}}{{5}}$ in the left (right) half, respectively, separated by a single column of local potential $J\sum_{y=1}^{13}\hat{n}_{(14,y)}$ in the middle. By numerically evaluating $\mathcal{C}(\vec{r},t)$ [Eq. (4)] and extracting the local Hall response $\sigma_{\rm H}(\vec{r})$ from a fit with Eq. (6), we find that the local Chern marker ${\sf Ch}(\vec{r})=2\pi\sigma_{\rm H}(\vec{r})$ takes values close to +1+1 (1-1) in the left (right) half of the system according to the sign of the magnetic flux, see Fig. 7, thus directly confirming its local nature.
Figure 7: Density and local Chern marker as obtained from the fit in Eq. (6) for a Harper-Hofstadter model with opposite flux $\pm\alpha=\nicefrac{{1}}{{5}}$ per plaquette in the left and right half separated by a single column of a local potential. As expected, we find that the local Chern marker ${\sf Ch}(\vec{r})$ changes sign as the reference site $\vec{r}$ is moved from one region to the other.
Including higher coefficients $c_{m}$. In order to improve the accuracy of our protocol, it is natural to include higher coefficients in the expansion of the correlator $\mathcal{C}(\vec{r},t)$ in Eq. (9),
@@
Acknowledgements.